### v1

# SLAM&Render: A Benchmark for the Intersection Between Neural Rendering, Gaussian Splatting and SLAM https://samuel-cerezo.github.io/SLAM&Render

Abstract
Models and methods originally developed for novel view synthesis and scene rendering, such as Neural Radiance Fields (NeRF) and Gaussian Splatting, are increasingly being adopted as representations in Simultaneous Localization and Mapping (SLAM). However, existing datasets fail to include the specific challenges of both fields, such as multi-modality and sequentiality in SLAM or generalization across viewpoints and illumination conditions in neural rendering. To bridge this gap, we introduce SLAM&Render, a novel dataset designed to benchmark methods in the intersection between SLAM and novel view rendering. It consists of 40 sequences with synchronized RGB, depth, IMU, robot kinematic data, and ground-truth pose streams. By releasing robot kinematic data, the dataset also enables the assessment of novel SLAM strategies when applied to robot manipulators. The dataset sequences span five different setups featuring consumer and industrial objects under four different lighting conditions, with separate training and test trajectories per scene, as well as object rearrangements. Our experimental results, obtained with several baselines from the literature, validate SLAM&Render as a relevant benchmark for this emerging research area.
I Introduction
Simultaneous Localization and Mapping, typically known by its acronym SLAM, is a key enabler for autonomous robot navigation and understanding. The goal is to simultaneously estimate the robot’s pose while creating a globally consistent representation or map of its surroundings, using only the data from the robot’s onboard sensors. After decades of intense research, several highly performing methods for geometric SLAM in rigid environments are available [1]. However, many research challenges remain [2]. SLAM differs from alternative approaches to 3D reconstruction (e.g., Structure from Motion [3]) by its online, real-time and multi-modal features, making it well-suited for robotic applications and requirements.
Novel view synthesis, seemingly distant from robot SLAM, has experienced a resurgence due to groundbreaking research in neural radiance fields, also known as NeRFs [4]. Nevertheless, it has been quickly adopted as a map representation in visual [5, 6] and LiDAR [7] SLAM pipelines. On the other hand, Gaussian Splatting [8], developed as a much faster alternative to NeRFs, has had an even bigger impact in visual SLAM [9, 10, 11, 12]. Refer to the recent survey by Tosi et al. [13] for a comprehensive overview of references and in-depth discussions on how these two paradigms are influencing the field of SLAM.
Our work is motivated by this emerging trend intersecting visual SLAM, neural rendering and Gaussian splatting. Specifically, we argue that there is a lack of public datasets and benchmarks that explicitly address this area of research. Therefore, we introduce SLAM&Render, a novel dataset for benchmarking methods at the intersection of the aforementioned fields. It includes 40 sequences with synchronized RGB, depth, IMU, robot kinematic data, intrinsic and extrinsic camera parameters and ground-truth data. The sequences cover five setups with consumer and industrial objects captured under varying lighting conditions and include separate training/test trajectories and object rearrangements.
All data is available online under the Creative Commons Attribution license (CC-BY 4.0) at the website. Additional information, example code and videos for visual inspection of the dataset is also included.
Figure 1: Illustration of the capture setup for our SLAM&Render dataset, recorded from a Intel RealSense at the end effector of a robotic arm that moves around a set of objects on a table. See the four different light conditions present in our dataset: a) natural, b) cold, c) warm, and d) dark
II Related Work
Year Dataset Use RGB Depth IMU Enc. GT Ill. Rearr. TS.
2014 DTU [14] NVS
2014 JIGSAWS [15] Motion Mod.
2015 YCB object and Model set [16] Robotic Manip.
2016 EuRoC [17] SLAM
2017 Tanks and Temples [18] NVS
2018 TUM VI [19] Odom. (✓)
2019 OpenLORIS-Scene [20] SLAM
2019 Replica [21] SLAM & NVS
2021 Mip-NeRF360 [22] NVS
2023 ScanNet++ [23] SLAM & NVS
2025 SLAM&Render SLAM & NVS
TABLE I: Overview of commonly used datasets for SLAM and NVS, including available sensor modalities (RGB, Depth, IMU, Encoders), ground truth camera pose (GT), illumination variations (Ill.), scene rearrangements (Rearr.), and time synchronization (TS).
Publicly available datasets have played a crucial role in advancing research in SLAM and Novel View Synthesis (NVS). In Table I, we provide an overview of the most relevant datasets that are instrumental to these fields of study.
II-A Datasets for SLAM
The EuRoC dataset [17] provides visual-inertial data from a MAV, comprising two batches: one collected in an industrial environment for evaluating visual SLAM algorithms, and the other in a room for 3D environment reconstruction. Several challenges such as varying image brightness and dynamic scenes, are included. The TUM VI dataset [19] is a visual-inertial dataset designed for evaluating odometry algorithms. It includes RGB images, photometric calibration, and synchronized IMU data. Ground-truth poses at the start and end of the sequences are provided. The dataset also includes evaluations of state-of-the-art methods. Depth images are not present in this collection. OpenLORIS-Scene dataset [20] consists of visual-inertial and odometry data to benchmark SLAM and scene understanding algorithms in everyday scenarios. For data collection a wheeled robot equipped with a Realsense D435i camera, encoders and a LiDAR sensor was used. In the context of synthetic datasets, the Replica dataset [21] contains 18 photorealistic indoor reconstructions at room and building scales. It features dense 3D meshes, HDR textures, and reflective surfaces. However, neither changes in illumination or rearrangements of objects within the same scene are included.
II-B Datasets for Novel View Synthesis.
Despite the popularity of NeRFs, most datasets used for evaluation sample frames along a single camera trajectory, often lacking the diversity and complexity of real-world settings. DTU [14] was introduced for evaluating multi-view stereo techniques using a robotic setup with a structured light scanner. It contains a wide variety of scenes with different object types and lighting conditions. Tanks and Temples [18] was developed for image-based 3D reconstruction. It was captured in realistic, non-laboratory settings using an industrial laser scanner for ground-truth. It includes both indoor and outdoor scenes. Recently, Mip-NeRF360 [22] introduced a dataset for high-quality novel view synthesis. It captured camera trajectories around a central object in indoor and outdoor scenarios with detailed backgrounds. Then, ScanNet++ [23] was introduced to benchmark 3D scene understanding methods. It combines 3D laser scans, DSLR images, and RGB-D sequences. It includes semantic annotations for diverse and ambiguous scenarios and is the first to use independent camera trajectories for NVS evaluation. However, it does not contain IMU, kinematic data or controlled illumination changes.
In the context of datasets containing kinematic data, the YCB Object and Model Set [16] provides object manipulation data with joint poses but does not include visual data. On the other hand, the JIGSAWS dataset [15] includes kinematic data from the da Vinci surgical robot and stereoscopic images of surgical tasks, providing a rich data source to study human-robot interaction. However, the dataset purpose makes it very domain-specific.
Unlike the existing dataset, the one presented in this work offers a comprehensive and diverse set of sequences designed to benchmark methods in the intersection between SLAM and novel view rendering. The dataset includes 40 sequences with synchronized streams of RGB images, depth, IMU, robot kinematic data, and ground-truth poses. The sequences cover five distinct setups with both consumer and industrial objects, captured under four lighting conditions. Each scene includes separate trajectories for training and testing, and variations in object arrangement to evaluate generalization. To the best of our knowledge, this is the first dataset containing multi-sensor and kinematic data collected in sequences characterized by varying lighting conditions, object rearrangements, occlusions, and time-synchronization (see Table I).
III Notation
Figure 2: Illustration of the objects included in SLAM&Render: a-b) supermarket goods, c) industrial goods, d)–h) object arrangements of the five setups.
III-A Convention for rigid transformations
Rigid transformations from a reference frame $A$ to another reference frame $B$ are defined as
$\bm{T}_{A}^{B}=\begin{bmatrix}\bm{R}_{A}^{B}&\bm{p}_{A}^{B}\\ \bm{0}_{1\times 3}&1\end{bmatrix}\in SE\left(3\right),$
where $\bm{R}_{A}^{B}\in SO\left(3\right)$ and $\bm{p}_{A}^{B}\in\mathbb{R}^{3}$ describe the orientation and the position of the frame $A$ with respect to the reference frame $B$, respectively. Such rigid transformation will transform a 3D point $\bm{p}^{B}\in\mathbb{R}^{3}$ ($\tilde{\bm{p}}^{B}$ in homogeneous coordinates) to the reference frame $A$ as $\tilde{\bm{p}}^{A}=\bm{T}_{A}^{B}\tilde{\bm{p}}^{B}$
III-B Forward kinematics
A manipulator consists of a sequence of rigid bodies (links) connected through kinematic joints. Let $\bm{x}\left(t\right)\in\mathbb{R}^{m}$ be the task variable to be controlled (e.g., end-effector pose) and $\bm{q}\left(t\right)\in\mathbb{R}^{n}$ the vector of the variables describing the configuration (i.e, joint positions) of a given $n$-DOFs robot, with $n\geq m$. The relation between $\bm{q}$ and $\bm{x}$ is given by the forward kinematics equation
$\bm{x}\left(t\right)=\bm{f}\left(\bm{q}\left(t\right)\right),$ (1)
where $\bm{f}\left(\cdot\right)$ is, in general, a nonlinear function allowing for the computation of the task variables starting from the robot configuration.
IV The SLAM&Render Dataset
The proposed dataset comprises 40 sequences that include both inertial and RGB-D data. These were recorded using an Intel RealSense D435i camera, mounted in an eye-to-hand configuration on a KUKA LBR iisy 3 R760 manipulator. Additionally, kinematic data were collected from the manipulator. In particular, the dataset contains the robot configurations $\bm{q}\left(t\right)$ along the motion and the corresponding flange poses computed using Eq. (1). An external Motion Capture System (MCS) was used to obtain the ground-truth camera poses. All sequences were recorded in an environment where the objects were placed on a black background, surrounded by black panels featuring pictures that display complex and colorful patterns (see Sec. IV-A). Moreover, the dataset contains train and test camera trajectories recorded under the following four different lighting conditions:
  • natural – Natural light of the environment
  • cold – Cold artificial light
  • warm – Warm artificial light
  • dark – No light.
[…]
  • setup-1 and setup-2 include a set of common supermarket goods including transparent objects
  • setup-3 includes a different set of common supermarket goods including only opaque objects
  • setup-4 contains industrial objects
  • setup-5 contains industrial objects in a cluttered arrangement.
We additionally release a collection of images capturing a chessboard calibration pattern to calibrate the intrinsic and extrinsic parameters of the camera. The calibration procedure, together with the estimation of extrinsic and intrinsic parameters, is detailed in Sec. IV-E.
In the website we provide a collection of tools for data association, retrieval from ROS2 bag files, alignment, and synchronization.
Figure 3: Our SLAM&Render dataset contains train and test camera trajectories. From top left to bottom right: a) 2D representation of train trajectory, b) 3D representation of train trajectory, c) 2D representation of test trajectory, and d) 3D representation of test trajectory. Blue and yellow points represent start and end points of the trajectories, respectively.
IV-A Data acquisition
All RGB-D data were recorded at the optimal depth resolution [^2] https://dev.intelrealsense.com/docs/tuning-depth-cameras-for-best-performance (848×480) and full frame rate (30 Hz) of the Intel RealSense D435i camera. The IMU data were acquired at a rate of 210 Hz, capturing synchronized readings from the gyroscope and accelerometer. All the dataset sequences were collected through a system based on Ubuntu 22.04 and ROS2 Humble. We used the software library Intel RealSense SDK 2.0333https://www.intelrealsense.com/sdk-2/ and its ROS2 wrapper to record the camera data. While the depth and color images are initially unaligned, as they are captured by two distinct sensors, we used the built-in feature of the Intel ROS2 wrapper to register the depth images to the color images. This is achieved by projecting the depth images to 3D and then back-projecting them into the color camera view. The joint angles of the robot manipulator were collected using ROS2 at a frequency of 25 Hz.
To record ground-truth data, we employed an OptiTrack Motion Capture System (MCS) composed of twelve PrimeX 13W cameras installed around the robot cell. These cameras track the position and orientation of passive markers at a frequency of 120 Hz. The markers are mounted on a circular structure between the camera and the robot flange (see Fig. 4). This structure was printed using an industrial-grade 3D printer, ensuring high accuracy while tracking the markers. Additionally, we have added two markers on the flange and one marker on the camera itself to enhance the asymmetry of the marker setup.
To minimize reflections from the surrounding environment, non-reflective black panels were installed around the perimeter of the robot cell. Pictures featuring complex and colorful patterns have been placed on the black panels not to oversimplify the novel view synthesis of NeRF-based algorithms. Further, the metal table supporting the robot manipulator was covered with a black cloth.
The cold and warm lighting setups were obtained thanks to a Neewer 660 PRO LED lamp (50 W).
IV-B Dataset structure
[…]
.1 train. .2 rgb/. .3 1739371345.939813614.png. .3 1739371346.193981361.png. .3 …. .2 depth/. .3 1739371345.939813614.png. .3 1739371346.193981361.png. .3 …. .2 robot_data/. .3 joint_positions.txt. .3 flange_poses.txt. .2 associations.txt. .2 imu.txt. .2 groundtruth_raw.csv. .2 groundtruth.txt.
These folders and files contain the following data:
  • rgb/: folder containing the timestamped color images (PNG format, 3 channels, 8 bits per channel);
  • depth/: folder containing the timestamped depth images (PNG format, 1 channel, 16 bits per channel, distance scaled by a factor of 1000). The depth images are already aligned with the RGB sensor;
  • associations.txt: file listing all the temporal associations between the RGB and depth images (format: rgb_timestamp rgb_file depth_timestamp depth_file);
  • imu.txt: file containing timestamped accelerometer and gyroscope readings (format: timestamp ax ay az gx gy gz);
  • groundtruth.txt: file containing the timestamped ground-truth of the camera poses w.r.t world reference frame, expressed as a sequence of quaternions and translation vectors (format: timestamp qx qy qz qw tx ty tz) (see Sec. IV-D);
  • groundtruth_raw.csv: file provided by the MCS containing ground-truth raw data of flange marker ring trajectory;
  • extrinsics.yaml: file containing the transformation matrices of the reference frames involved in the data collection process (see Sec. IV-C);
  • instrinsics.yaml: file containing the intrinsic calibration parameters of the camera (see Sec. IV-E);
  • robot_data/flange_poses.txt: file containing the timestamped poses of the robot flange. Each pose is expressed w.r.t. the robot base as a quaternion and translation vector (format: timestamp qx qy qz qw tx ty tz);
  • robot_data/joint_positions.txt: file reporting the timestamped angles of each robot joint (format: timestamp q1 q2 q3 q4 q5 q6).
The accelerometer data are provided in $m/s^{2}$, the gyroscope data in $rad/s$, and the ground-truth and flange positions in $mm$. All sequences are also available as ROS2 bag files.
IV-C Reference frames
The reference frames shown in Fig. 4 were used in the data collection process to align the data coming from the camera, robot, and MCS. These reference frames are denoted as follows:
  • robot_flange: defined at the center of the real robot flange;
  • flange_marker_ring: used by the MCS to locate the robot flange;
  • rgb_sensor: optical center of the RGB sensor;
  • imu: describing the IMU sensor pose;
  • base_marker_ring: used by the MCS to locate the robot base;
  • robot_base: defined at the center of the real robot base.
Figure 4: Illustration of the reference frames involved in the data collection.
IV-D Motion Capture System (MCS)
[…]
Figure 5: Validation of time-synchronized data comparing the angular velocity estimated from the MCS and from the IMU gyroscope measurements.
Synchronization. The data from the MCS are initially unsynchronized. To ensure temporal alignment, for each sequence we retrieved the capture starting time provided by OptiTrack. Consequently, the ground-truth file we provide includes the corresponding Unix Epoch timestamp for each pose, calculated using using the capture starting time. Following the approach presented in [25], the synchronization was validated by comparing the angular velocity estimated from the MCS with the gyroscope measurements from the IMU using cross-correlation (see Fig. 5). The maximum at zero offset confirms correct alignment.
IV-E Camera calibration
To estimate the hand-eye transformation (i.e., the transformation from the robot flange to the camera projection plane), more than 200 images containing the calibration plate were captured, and the corresponding robot pose for each image was saved. Subsequently, the calibration method presented in [26] was applied. The calibration plate (a chessboard with squares of $30\times 30~{}mm$) was detected by applying a custom detection method (see Fig. 6). The intrinsics of the RealSense D435i camera were refined from the factory calibration ones.
Figure 6: Corners and fiducial markers detected on the calibration plate.
For a more accurate calibration, not only the extrinsic parameters (hand-eye transformation) were estimated, but also the intrinsic parameters. Our experiments have proven that the use of the estimated intrinsic parameters resulted in more accurate extrinsic parameters. The intrinsic parameters of the camera (focal length (fx/fy), principal point (cx/cy) and camera distortion coefficients) are shown in Table II. The extrinsic parameters are shown in the Table III.
Parameters fx fy cx cy distortion coefficients
factory 605.2 605.1 425.7 246.0 0 0 0 0 0
estimated 599.6 599.2 426.6 246.3 0.1 -0.4 0 0 0.3
TABLE II: Intrinsic parameters of the camera in [px] used in our dataset.
Used intrinsics tx ty tz qx qy qz qw
factory 0.5895 32.5419 46.0707 0.0060 0.0041 -0.7035 0.7106
estimated -0.2852 32.7894 46.1513 0.0063 0.0054 -0.7037 0.7104
TABLE III: Extrinsic parameters (hand-eye transformation) of the camera, represented as a translation [$mm$] and quaternion.
Reproj. error Hand-eye
[px] Trans. [$mm$] Rot. [$deg.$] Rel. transf. err. [$a.u.$]
Mean 0.36 2.38 0.33 2.90
Std. dev. 0.04 0.41 0.08 0.66
Std. error 0.01 0.13 0.02 0.21
All images (estim. intrinsics) 0.38 2.31 0.31 2.69
All images (factory intrinsics) 0.61 2.37 0.33 2.77
TABLE IV: Mean, standard deviation and standard error of calibration metrics. The last two rows show the results using all images with the estimated and factory intrinsic parameters.
To compare the accuracy of the calibration using both factory and estimated intrinsic parameters, the set of calibration images was divided into several batches (10 batches with 20 images in each), and the same calibration process was applied to each batch. The accuracy of the calibration was assessed using the metrics proposed in [27], which include translation error (Trans. [$mm$]), rotation error (Rot. [$deg.$]), and relative transformation error (Rel. transf. error [$a.u.$]).
Table IV shows the mean, standard deviation, and standard error of the calibration metrics. The accuracy of the hand-eye transformation, calculated using the estimated intrinsic parameters, showed more precise results, with reduced reprojection, translation, rotation, and relative transformation errors.
V Experiments
V-A Metrics
In this work, we report the Root Mean Square Error of the Absolute Trajectory Error (ATE RMSE $\downarrow$) to assess the camera tracking accuracy. To evaluate novel view synthesis, we follow the evaluation protocol presented in [8] and report standard photometric rendering metrics such as Peak Signal-to-Noise Ratio (PSNR $\uparrow$), Structural Similarity Index Measure (SSIM $\uparrow$), and Learned Perceptual Image Patch Similarity (LPIPS $\downarrow$). PSNR captures pixel-level differences, SSIM reflects perceptual similarity in terms of structure, luminance, and contrast [28]. LPIPS evaluates visual fidelity based on deep feature representations [29].
V-B Results
To evaluate the dataset’s suitability for benchmarking methods at the intersection of SLAM and novel view synthesis, we present results from state-of-the-art techniques.
Novel View Synthesis. To highlight the usefulness of independent test camera trajectories, we evaluated the performance of two state-of-the-art baselines, Gaussian Splatting [8] and FeatSplat [30], this last one specifically designed to reduce viewpoint overfitting. In this experiment, setup-1, setup-3, and setup-4 were used under natural, warm, and dark lighting conditions. Baselines were initialized with random point-clouds and a black background, using 4 Spherical Harmonics bands for Gaussian Splatting, and 32-dimensional latent vectors and 1 hidden layer of 64 neurons MLP for FeatSplat. They were optimized for a total of 21k iterations, and we report metrics both at 7k and 21k iterations. The baselines were optimized on the respective train trajectories of each setup, skipping one every ten frames. We evaluate both with the common approach of using skipped frames from the train trajectory [4, 8], Table V, and using the independently recorded test trajectory of each setup, Table VI.
natural warm dark
PSNR LPIPS SSIM PSNR LPIPS SSIM PSNR LPIPS SSIM
gsplat-7k 17.32 0.445 0.694 17.67 0.497 0.609 23.77 0.464 0.678
fsplat-7k 17.40 0.445 0.700 17.75 0.499 0.610 23.43 0.467 0.687
gsplat-21k 18.43 0.411 0.724 18.75 0.468 0.637 24.99 0.443 0.708
fsplat-21k 19.45 0.406 0.743 20.01 0.461 0.656 25.13 0.444 0.714
TABLE V: Classic evaluation of Gaussian Splatting (gsplat) and FeatSplat (fsplat) reporting PSNR[$\uparrow$], LPIPS[$\downarrow$], and SSIM[$\uparrow$].
natural warm dark
PSNR LPIPS SSIM PSNR LPIPS SSIM PSNR LPIPS SSIM
gsplat-7k 14.53 0.493 0.608 15.45 0.543 0.527 21.82 0.499 0.596
fsplat-7k 14.73 0.488 0.623 15.42 0.541 0.540 21.73 0.498 0.630
gsplat-21k 14.42 0.478 0.606 15.35 0.531 0.524 21.53 0.492 0.600
fsplat-21k 14.71 0.478 0.614 15.29 0.532 0.535 21.91 0.490 0.627
TABLE VI: Evaluation on independent test trajectory of Gaussian Splatting (gsplat) and FeatSplat (fsplat) reporting PSNR[$\uparrow$], LPIPS[$\downarrow$], and SSIM[$\uparrow$]
Figure 7: Comparison of Novel View Synthesis for validation and test viewpoints. Despite both models perform well when evaluated on validation viewpoints of the training trajectory, the independent test trajectory reveals overfitting.
The common evaluation indicates an improvement in rendering performance, for both baselines, through the complete optimization process, see Table V. Both baselines achieve their best performance under dark lighting conditions, which is probably due to using black as initial background color. Nevertheless, the test trajectory at SLAM&Render reveals that both baselines overfit to the training trajectory. Note how the rendering quality does not improve beyond the 7k iteration and even slightly declines in Table Table VI. Qualitative analysis in Fig. 7 also showcases the contrast between the high quality rendering of validation viewpoints and the artifacts generated in test viewpoints. This result validates the need for datasets with independent trajectories to evaluate novel view synthesis methods and shows that the classic approach of sampling evaluation frames from the training trajectory does not inform about trajectory overfitting.
Setup Ill. Traj. MonoGS Kinematics MonoGS (Kinem. seed)
setup-1 natural train 3.3 0.5 2.0
natural test 3.9 0.4 2.1
cold test 2.4 0.3 2.2
warm test 2.1 0.3 1.4
dark test 3.1 0.4 1.9
setup-3 natural train 2.6 0.5 1.1
natural test 3.7 0.4 1.7
dark train 0.6 0.5 0.6
dark test 3.8 0.3 1.6
setup-4 natural train 0.9 0.5 0.7
cold train 2.4 0.5 1.0
warm train 1.6 0.5 0.8
dark train 1.0 0.5 0.9
TABLE VII: ATE RMSE [$\downarrow cm$] on different sequences running three algorithm configurations: MonoGS, Kinematics and MonoGS with Kinematic data as seed
Camera tracking. To illustrate the advantage on using kinematic data, three approaches based on Gaussian Splatting SLAM [9] (MonoGS) were evaluated. In the first experiment, we used the default configuration for MonoGS. In the second experiment we incorporated the robot kinematic data directly as the camera motion. This approach is referred to as the Kinematics experiment. In the third experiment, robot kinematrics were incorporated as an initial seed for camera pose estimation within MonoGS. This is referred to as Kinem. seed experiment. Following the evaluation conducted by the authors of MonoGS in [9], the ATE RMSE for keyframes is reported in Table VII. For this experiment, setup-1, setup-3, and setup-4 were used under different light conditions while performing train and test trajectories. In all cases, it is evident that incorporating kinematic data improves camera tracking performance. However, the results also show that the naive approach of using robot kinematics as the initial seed of MonoGS decreases its accuracy. This result motivates the use of our SLAM&Render dataset for research on multimodal fusion within NeRF and Gaussian Splatting SLAM. Fig. 8 illustrates the camera tracking results for setup-4 under cold lighting condition. The trajectory using MonoGS is shown in Fig. 8a, while Fig. 8b shows the result using kinematic data is used as initial seed. Incorporating kinematic data allows drift correction in camera tracking.
Figure 8: Camera tracking result for setup-4 with cold light using a) MonoGS and b) MonoGS with kinematic data as initial seed. Dashed line represent the ground-truth while colored line represent the camera trajectory. Incorporating kinematic data allows drift correction in camera tracking.
VI Conclusions
In this work we introduced SLAM&Render, a dataset for benchmarking methods that combine SLAM and novel view rendering. Our data includes 40 sequences of synchronized RGB and depth images, IMU readings, robot kinematic data, intrinsic and extrinsic camera parameters, and ground-truth pose data across diverse scenes and conditions. SLAM&Render addresses research challenges such as generalization across trajectories and robustness to scene and illumination changes. It aims to drive future research in learning-based and hybrid methods requiring accurate localization and multi-sensor data fusion, ultimately enhancing the deployment of systems that combine perception, mapping, and rendering.
VII Acknowledgments
The authors would like to thank the Technology & Innovation Center of KUKA Deutschland GmbH (Augsburg, Germany) for their support and resources provided during the course of this research.